# MTMamba++: Enhancing Multi-Task Dense Scene Understanding via Mamba-Based Decoders

### v1

@@ Abstract
Multi-task dense scene understanding, which trains a model for multiple dense prediction tasks, has a wide range of application scenarios. Capturing long-range dependency and enhancing cross-task interactions are crucial to multi-task dense prediction. In this paper, we propose MTMamba++, a novel architecture for multi-task scene understanding featuring with a Mamba-based decoder. It contains two types of core blocks: self-task Mamba (STM) block and cross-task Mamba (CTM) block. STM handles long-range dependency by leveraging state-space models, while CTM explicitly models task interactions to facilitate information exchange across tasks. We design two types of CTM block, namely F-CTM and S-CTM, to enhance cross-task interaction from feature and semantic perspectives, respectively. Experiments on NYUDv2, PASCAL-Context, and Cityscapes datasets demonstrate the superior performance of MTMamba++ over CNN-based and Transformer-based methods. The code is available at https://github.com/EnVision-Research/MTMamba.
@@ 1 Introduction
Multi-task dense scene understanding has a variety of practical applications in computer vision [1], such as autonomous driving [2], healthcare [3], and robotics [4]. This problem is challenging because it requires training a model to simultaneously handle multiple dense prediction tasks, such as semantic segmentation, monocular depth estimation, surface normal estimation, and object boundary detection.
The widely adopted architecture for multi-task dense scene understanding is based on the encoder-decoder framework, which features a common encoder for extracting task-shared features and multiple decoders for generating task-specific predictions [1]. This framework is very general and many variants [5, 6, 7, 8, 9, 10] have been proposed to improve its performance. One promising approach is the decoder-focused method [1], which aims to enhance cross-task interaction in task-specific decoders by well-designed fusion modules. For example, derived from convolutional neural networks (CNNs), PAD-Net [5] and MTI-Net [6] use a multi-modal distillation module to enhance information exchange across tasks in the decoder and achieve better performance than the conventional encoder-decoder framework.
Since the convolution operation mainly captures local features and may struggle with modeling long-range dependencies [11, 12], recent methods [7, 8, 9, 10] adopt Transformer-based decoders with attention-based fusion modules. These methods leverage the attention mechanism [13] to capture global context information, resulting in better performance than CNN-based methods. These works demonstrate that enhancing cross-task correlation and modeling long-range spatial relationships are crucial to multi-task dense prediction.
Recently, Mamba [14], a new network architecture derived from state space models (SSMs) [15, 16], has demonstrated better capacity in long-range dependencies modeling and superior performance than Transformers in various domains, including language modeling [14, 17, 18], graph reasoning [19, 20], medical images analysis [21, 22], and image generation [23, 24]. However, existing works on Mamba are limited to single-task learning scenarios. Using Mamba to solve multi-task problems is still unexplored since it needs to model the cross-task interaction, which is challenging.
To fill these gaps, in this paper, we propose MTMamba++, a novel multi-task architecture featuring a Mamba-based decoder. As the overall framework shown in Figure 1, MTMamba++ is a decoder-focused method and the decoder consists of three stages. These three stages share a similar structure and contain three types of blocks: ECR, STM, and CTM blocks. Specifically, the ECR (expand, concatenate, and reduce) block is responsible for upscaling the task feature and fusing it with the high-level feature from the encoder. It allows each decoder stage to incrementally increase the spatial resolution of the feature maps, which plays an important role in precise pixel-level predictions. The core components of the decoder are two Mamba-based blocks, i.e., the STM and CTM blocks, as shown in Figures 2(b) and 3, respectively. The self-task Mamba (STM) block, inspired by [25], effectively captures global context information for each task by the SSM mechanism [14]. The cross-task Mamba (CTM) block is designed to enhance each task’s features by exchanging knowledge among various tasks. We develop two variants of the CTM block, called F-CTM and S-CTM. F-CTM promotes task interaction at the feature level, where each task can dynamically integrate the cross-task representation with its own task features through a task-specific and input-dependent gate. S-CTM, on the other hand, delves into the intricate relationships between tasks at a semantic level by employing a novel cross SSM, which allows modeling the relationship between two sequences based on the SSM mechanism. By integrating STM and CTM blocks in the decoder, MTMamba++ not only enhances cross-task interaction but also effectively handles long-range dependency. We evaluate MTMamba++ on three standard multi-task dense prediction benchmark datasets, namely NYUDv2 [26], PASCAL-Context [27], and Cityscapes [28]. Quantitative results demonstrate that MTMamba++ significantly surpasses both CNN-based and Transformer-based methods. Additionally, qualitative studies show that MTMamba++ generates superior visual results with greater accuracy in detail compared to state-of-the-art Transformer-based methods.
@@
  • We propose MTMamba++, a novel multi-task architecture based on state space models (SSMs) for multi-task dense scene understanding. It contains a novel Mamba-based decoder, which effectively models long-range spatial relationships and achieves cross-task correlation;
@@
  • We evaluate MTMamba++ on three benchmark datasets, including NYUDv2, PASCAL-Context, and Cityscapes. Quantitative results demonstrate the superiority of MTMamba++ on multi-task dense prediction over previous CNN-based and Transformer-based methods. Qualitative evaluations show that MTMamba++ generates precise predictions.
@@ 2.1 Multi-Task Learning
Multi-task learning (MTL) is a learning paradigm that aims to jointly learn multiple related tasks using a single model [31]. Current MTL research mainly focuses on multi-objective optimization [32, 33, 34, 35, 36, 37] and network architecture design [7, 9, 10, 5, 6, 8]. In multi-task visual scene understanding, most existing works focus on designing architecture [1], especially developing specific modules in the decoder to facilitate knowledge exchange among different tasks. For instance, based on CNN, Xu et al. [5] introduce PAD-Net, which integrates an effective multi-modal distillation module aimed at enhancing information exchange among various tasks within the decoder. MTI-Net [6] is a complex multi-scale and multi-task CNN architecture that facilitates information distillation across various feature scales. As the convolution operation only captures local features [11], recent approaches [7, 9, 10, 8] develop Transformer-based decoders to grasp global context by attention mechanism [13]. For example, InvPT [7] is a Transformer-based multi-task architecture that employs an effective UP-Transformer block for multi-task feature interaction at different feature scales. MQTransformer [8] uses a cross-task query attention module in the decoder to enable effective task association and information communication.
@@ 2.2 State Space Models
Recently, Gu and Dao [14] propose a new SSM-based architecture termed Mamba, which incorporates a new SSM called S6. This SSM is an input-dependent selection mechanism derived from S4. Mamba has demonstrated superior performance over Transformers on various benchmarks, such as image classification [25, 45], image segmentation [22], and graph prediction [19]. Different from existing research efforts on Mamba, which mainly focus on single-task settings, in this paper, we consider a more challenging multi-task setting and propose a novel cross-task Mamba module to capture inter-task dependence.
@@ 3.4 ECR Block
The ECR (expand, concatenate, and reduce) block is responsible for upsampling the feature and aggregating it with the encoder’s feature. As illustrated in Figure 2(a), it contains three steps. For an input feature, ECR block first 2×2$\times$2 × upsamples the feature resolution and 2×2$\times$2 × reduces the channel number by a linear layer and the rearrange operation. Then, the feature is fused with the high-level feature from the encoder through skip connections. Fusing these features is crucial for compensating the loss of spatial information that occurs due to downsampling in the encoder. Finally, a $1\times 1$ convolutional layer is used to reduce the channel number. Consequently, the ECR block facilitates the efficient recovery of high-resolution details, which is essential for dense prediction tasks that require precise spatial information.
@@ 3.7.2 LiteHead
In DenseHead, upsampling is performed first, which can lead to a significant computational cost. Hence, we introduce a more simple, lightweight, and effective head architecture, called LiteHead. Specifically, it consists of a $3\times 3$ convolutional layer, followed by a batch normalization layer, a ReLU activation function, and a final linear layer that projects the feature channels onto the task’s output dimensions. Subsequently, the feature is simply interpolated to align with the input resolution and then used as the output. Thus, LiteHead is much more computationally efficient than DenseHead. Note that since each task has its own head, the overall computational cost reduction is linearly related to the number of tasks.
@@ 4.1.1 Datasets
Following [7, 9, 10], we conduct experiments on three multi-task dense prediction benchmark datasets: (i) NYUDv2 [26] contains a number of indoor scenes, including 795 training images and 654 testing images. It consists of four tasks: $40$-class semantic segmentation (Semseg), monocular depth estimation (Depth), surface normal estimation (Normal), and object boundary detection (Boundary). (ii) PASCAL-Context [27], originated from the PASCAL dataset [47], includes both indoor and outdoor scenes and provides pixel-wise labels for tasks like semantic segmentation, human parsing (Parsing), and object boundary detection, with additional labels for surface normal estimation and saliency detection tasks generated by [48]. It contains 4,998 training images and 5,105 testing images. (iii) Cityscapes [28] is an urban scene understanding dataset. It has two tasks (19-class semantic segmentation and disparity estimation) with 2,975 training and 500 testing images.
@@ 4.1.2 Implementation Details
We use the Swin-Large Transformer [30] pretrained on the ImageNet-22K dataset [49] as the encoder. The expansion factor $\alpha$ is set to $2$ in both STM and CTM blocks. Following [7, 9, 10], we resize the input images of NYUDv2, PASCAL-Context, and Cityscapes datasets as $448\times 576$, $512\times 512$, and $512\times 1024$, respectively, and use the same data augmentations including random color jittering, random cropping, random scaling, and random horizontal flipping. The $\ell_{1}$ loss is used for depth estimation and surface normal estimation tasks, while the cross-entropy loss is for other tasks. The proposed model is trained with a batch size of 4 for 40,000 iterations. The AdamW optimizer [50] with a weight decay of $1\times 10^{-6}$ and the polynomial learning rate scheduler are used for all three datasets. The learning rate is set to $2\times 10^{-5}$, $8\times 10^{-5}$, and $1\times 10^{-4}$ for NYUDv2, PASCAL-Context, and Cityscapes datasets, respectively.
@@
TABLE II: Comparison with state-of-the-art methods on the Cityscapes dataset. ($\uparrow$($\downarrow$)↑ ( ↓ ) indicates that a higher (lower) result corresponds to better performance. The best and second best results are highlighted in bold and underline, respectively.
@@ 4.1.3 Evaluation Metrics
where $T$ is the number of tasks, $M_{t}^{\mathcal{A}}$ is the metric value of method $\mathcal{A}$ on task $t$, and $s_{t}$ is 00 if a larger value indicates better performance for task $t$, and $1$ otherwise.
@@ 4.2 Comparison with State-of-the-art Methods
We compare the proposed MTMamba++ method with two types of MTL methods: (i) CNN-based methods, including Cross-Stitch [51], PAP [52], PSD [53], PAD-Net [5], MTI-Net [6], ATRC [54], and ASTMT [48]; and (ii) Transformer-based methods, including InvPT [7], TaskPrompter [10], InvPT++ [9], and MQTransformer [8].
Table I provides the results on NYUDv2 and PASCAL-Context datasets. As can be seen, Mamba-based methods largely outperform CNN-based and Transformer-based methods, especially MTMamba++, which achieves the best performance in all four tasks of NYUDv2. Specifically, MTMamba++ significantly outperforms the previous state-of-the-art Transformer-based method (i.e., TaskPrompter) by +1.71 (mIoU) and +1.20 (odsF) in semantic segmentation and object boundary detection tasks, respectively, which demonstrates the superiority of MTMamba++. Moreover, MTMamba++ performs better than MTMamba, showing the effectiveness of S-CTM and LiteHead. The results on PASCAL-Context also show the clear superiority of Mamba-based methods. Notably, MTMamba++ surpasses the Transformer-based approaches on all tasks except the normal prediction task, which is also comparable. Specifically, compared with the best Transformer-based approaches, MTMamba++ achieves improvements of +1.05 (mIoU), +3.75 (mIoU), and +4.40 (odsF) in semantic segmentation, human parsing, and object boundary detection tasks, respectively, which demonstrates the superior efficacy of MTMamba++. Moreover, compared with MTMamba, MTMamba++ outperforms on three of five tasks and achieves comparable performance on the other two tasks, demonstrating the effectiveness of MTMamba++ again.
Table II shows the results on the Cityscapes dataset. We can see that Mamba-based methods perform largely better than the previous CNN-based and Transformer-based approaches on both two tasks. Moreover, MTMamba++ archives the best performance. Notably, MTMamba++ outperforms the current state-of-the-art TaksPrompter by +6.72 (mIoU) in the semantic segmentation task, demonstrating that MTMamba++ is more effective. Besides, MTMamba++ performs better than MTMamba, which shows the effectiveness of S-CTM and LiteHead.
TABLE III: Effectiveness of each core component on NYUDv2. “Multi-task” denotes an MTL model where each task only uses two standard Swin Transformer blocks [30] after the ECR block in each decoder stage. “Single-task” is the single-task counterpart of “Multi-task”. #7 is the default configuration of MTMamba++.
@@
The qualitative comparisons with state-of-the-art methods (i.e., InvPT [7], TaskPrompter [10], and MTMamba [29]) on NYUDv2, PASCAL-Context, and Cityscapes datasets are shown in Figures 5, 6, and 7, demonstrating that MTMmaba++ provides more precise predictions and details.

### v2

@@ Abstract
Multi-task dense scene understanding, which trains a model for multiple dense prediction tasks, has a wide range of application scenarios. Capturing long-range dependency and enhancing cross-task interactions are crucial to multi-task dense prediction. In this paper, we propose MTMamba++, a novel architecture for multi-task scene understanding featuring with a Mamba-based decoder. It contains two types of core blocks: self-task Mamba (STM) block and cross-task Mamba (CTM) block. STM handles long-range dependency by leveraging state-space models, while CTM explicitly models task interactions to facilitate information exchange across tasks. We design two types of CTM block, namely F-CTM and S-CTM, to enhance cross-task interaction from feature and semantic perspectives, respectively. Extensive experiments on NYUDv2, PASCAL-Context, and Cityscapes datasets demonstrate the superior performance of MTMamba++ over CNN-based, Transformer-based, and diffusion-based methods while maintaining high computational efficiency. The code is available at https://github.com/EnVision-Research/MTMamba.
@@ 1 Introduction
Multi-task dense scene understanding, which trains a single model to simultaneously handle multiple pixel-wise prediction tasks (e.g., semantic segmentation, depth estimation, surface normal estimation, and object boundary detection), has become increasingly important in many computer vision applications [1], such as autonomous driving [2], healthcare [3], and robotics [4]. The success of multi-task dense prediction hinges on addressing two fundamental challenges: (i) modeling long-range spatial relationships to capture global context information, which is essential for pixel-wise prediction tasks; (ii) enhancing cross-task interactions to facilitate knowledge sharing among different tasks, which is crucial to multi-task learning.
Existing multi-task dense prediction approaches can be broadly categorized by their architectural design. CNN-based methods [5, 6] employ convolutional operations in decoders for task-specific predictions but primarily capture local features, struggling with modeling long-range dependencies and global context understanding [7, 8]. Transformer-based methods [9, 10, 11, 12, 13, 14] employ attention mechanisms [15] to better capture global context and demonstrate improved performance. However, they suffer from quadratic computational complexity with respect to sequence length [16, 17], making them computationally prohibitive for high-resolution dense prediction tasks.
To address these limitations, we propose MTMamba++, a novel Mamba-based architecture that achieves effective and efficient multi-task dense scene understanding. MTMamba++ introduces two key components based on state space models (SSMs) [18, 19] in the decoder: (i) The self-task Mamba (STM) block, inspired by [20], captures global context information for each task by leveraging the long-range modeling capabilities of SSMs with linear computational complexity; (ii) The cross-task Mamba (CTM) block enables effective knowledge exchange across tasks through two variants: F-CTM for feature-level interaction and S-CTM for semantic-level interaction. The S-CTM introduces a novel cross SSM (CSSM) mechanism that models relationships between task-specific and shared feature sequences, providing more effective task interaction than simple feature fusion approaches used in F-CTM.
As the overall framework shown in Figure 1, MTMamba++ features a three-stage Mamba-based decoder that progressively refines multi-task predictions. Each stage contains an ECR (expand, concatenate, and reduce) block that upscales features and fuses them with encoder features, followed by STM and CTM blocks for task-specific learning and cross-task interaction. This design effectively captures long-range dependencies and enhances cross-task interaction while maintaining computational efficiency.
We evaluate MTMamba++ on three standard multi-task dense prediction benchmark datasets, namely NYUDv2 [21], PASCAL-Context [22], and Cityscapes [23]. Quantitative results demonstrate that MTMamba++ significantly surpasses previous methods, including CNN-based, Transformer-based, and diffusion-based appoarch. Moreover, comprehensive efficiency analysis shows that MTMamba++ achieves state-of-the-art performance while maintaining high computational efficiency. Notably, our experiments demonstrate that SSM-based architectures are more effective and efficient than attention-based for multi-task dense prediction tasks. Additionally, qualitative studies show that MTMamba++ generates superior visual results with greater accuracy in detail, sharper boundaries, and more accurate detection in small objects compared to existing approaches.
@@
  • We propose MTMamba++, a novel multi-task architecture based on state space models (SSMs) for multi-task dense scene understanding. It contains a novel Mamba-based decoder, effectively modeling long-range spatial relationships and achieving cross-task correlation;
@@
  • We evaluate MTMamba++ on three benchmark datasets, including NYUDv2, PASCAL-Context, and Cityscapes. Quantitative results demonstrate the superiority of MTMamba++ on multi-task dense prediction over previous methods while maintaining high computational efficiency. Qualitative evaluations show that MTMamba++ generates precise predictions.
@@ 2.1 Multi-Task Learning
Multi-task learning (MTL) is a learning paradigm that aims to jointly learn multiple related tasks using a single model [26, 27]. Current MTL research mainly focuses on multi-objective optimization [28, 29, 30, 31, 32, 33] and network architecture design [9, 11, 12, 5, 6, 10, 13, 14, 34, 35]. In multi-task visual scene understanding, most existing works focus on designing architecture [1], especially developing specific modules in the decoder to facilitate knowledge exchange among different tasks. For instance, based on CNN, Xu et al. [5] introduce PAD-Net, which integrates an effective multi-modal distillation module aimed at enhancing information exchange among various tasks within the decoder. MTI-Net [6] is a complex multi-scale and multi-task CNN architecture that facilitates information distillation across various feature scales. As the convolution operation only captures local features [7], recent approaches [9, 11, 12, 10, 13, 14] develop Transformer-based decoders to grasp global context by attention mechanism [15]. For example, InvPT [9] is a Transformer-based multi-task architecture that employs an effective UP-Transformer block for multi-task feature interaction at different feature scales. MQTransformer [10] uses a cross-task query attention module in the decoder to enable effective task association and information communication.
@@ 2.2 State Space Models
Recently, Gu and Dao [36] propose a new SSM-based architecture termed Mamba, which incorporates a new SSM called S6. This SSM is an input-dependent selection mechanism derived from S4. Mamba has demonstrated superior performance over Transformers on various benchmarks, such as language modeling [36, 42, 43], graph reasoning [44, 45], medical image analysis [46, 47], and image classification [20, 48]. Different from existing research efforts on Mamba, which mainly focus on single-task settings, in this paper, we consider a more challenging multi-task setting and propose a novel cross-task Mamba module to capture inter-task dependence.
@@ 3.4 ECR Block
The ECR (expand, concatenate, and reduce) block is responsible for upsampling the feature and aggregating it with the encoder’s feature. As illustrated in Figure 2(a), it contains three steps. For an input feature, ECR block first 2×2$\times$2 × upsamples the feature resolution and 2×2$\times$2 × reduces the channel number by a linear layer and the rearrange operation. Then, the feature is fused with the high-level feature from the encoder through skip connections. Fusing these features is crucial for compensating the loss of spatial information that occurs due to downsampling in the encoder. Finally, a $\times 1$ 1 convolutional layer is used to reduce the channel number. Consequently, the ECR block facilitates the efficient recovery of high-resolution details, which is essential for dense prediction tasks that require precise spatial information.
@@ 3.7.2 LiteHead
In DenseHead, upsampling is performed first, which can lead to a significant computational cost. Hence, we introduce a more simple, lightweight, and effective head architecture, called LiteHead. Specifically, it consists of a $\times 3$ 3 convolutional layer, followed by a batch normalization layer, a ReLU activation function, and a final linear layer that projects the feature channels onto the task’s output dimensions. Subsequently, the feature is simply interpolated to align with the input resolution and then used as the output. Thus, LiteHead is much more computationally efficient than DenseHead. Note that since each task has its own head, the overall computational cost reduction is linearly related to the number of tasks.
@@ 4.1.1 Datasets
Following [9, 11, 12], we conduct experiments on three multi-task dense prediction benchmark datasets: (i) NYUDv2 [21] contains a number of indoor scenes, including 795 training images and 654 testing images. It consists of four tasks: 404040-class semantic segmentation (Semseg), monocular depth estimation (Depth), surface normal estimation (Normal), and object boundary detection (Boundary). (ii) PASCAL-Context [22], originated from the PASCAL dataset [50], includes both indoor and outdoor scenes and provides pixel-wise labels for tasks like semantic segmentation, human parsing (Parsing), and object boundary detection, with additional labels for surface normal estimation and saliency detection tasks generated by [51]. It contains 4,998 training images and 5,105 testing images. (iii) Cityscapes [23] is an urban scene understanding dataset. It has two tasks (19-class semantic segmentation and disparity estimation) with 2,975 training and 500 testing images.
@@ 4.1.2 Implementation Details
We use the Swin-Large Transformer [25] pretrained on the ImageNet-22K dataset [52] as the encoder. The expansion factor $\alpha$ is set to 222 in both STM and CTM blocks. Following [9, 11, 12], we resize the input images of NYUDv2, PASCAL-Context, and Cityscapes datasets as 448×576448$\times$ 576448 × 576, $\times 512$ 512, and 512×1024512$\times$ 1024512 × 1024, respectively, and use the same data augmentations including random color jittering, random cropping, random scaling, and random horizontal flipping. The $\ell_{1}$ loss is used for depth estimation and surface normal estimation tasks, while the cross-entropy loss is for other tasks. The proposed model is trained with a batch size of 4 for 40,000 iterations. The AdamW optimizer [53] with a weight decay of $1\times 10^{-6}$ and the polynomial learning rate scheduler are used for all three datasets. The learning rate is set to $2\times 10^{-5}$, $8\times 10^{-5}$, and $1\times 10^{-4}$ for NYUDv2, PASCAL-Context, and Cityscapes datasets, respectively.
@@
TSP-Transformer [13] 55.39 0.4961 18.44 77.50
MLoRE [14] 55.96 0.5076 18.33 78.43
Diffusion-based decoder
TaskDiffusion [34] 55.65 0.5020 18.43 78.64
@@
TSP-Transformer [13] 81.48 70.64 84.86 13.69 74.80
MLoRE [14] 81.41 70.52 84.90 13.51 75.42
Diffusion-based decoder
TaskDiffusion [34] 81.21 69.62 84.94 13.55 74.89
@@
TABLE II: Comparison with state-of-the-art methods on the Cityscapes dataset. ($\uparrow$($\downarrow$)↑ ( ↓ ) indicates that a higher (lower) result corresponds to better performance. The best and second-best results are highlighted in bold and underline, respectively.
@@ 4.1.3 Evaluation Metrics
where $T$ is the number of tasks, $M_{t}^{\mathcal{A}}$ is the metric value of method $\mathcal{A}$ on task $t$, and $s_{t}$ is 0 if a larger value indicates better performance for task $t$, and 111 otherwise.
@@ 4.2 Comparison with State-of-the-art Methods
We compare the proposed MTMamba++ method with two types of MTL methods: (i) CNN-based methods, including Cross-Stitch [54], PAP [55], PSD [56], PAD-Net [5], MTI-Net [6], ATRC [57], and ASTMT [51]; (ii) Transformer-based methods, including InvPT [9], TaskPrompter [12], InvPT++ [11], MQTransformer [10], TSP-Transformer [13], and MLoRE [14]; and (iii) Diffusion-based method TaskDiffusion [34].
Table I provides the results on NYUDv2 and PASCAL-Context datasets. As can be seen, MTMamba++ largely outperforms CNN-based, Transformer-based, and Diffusion-based methods, especially achieving the best performance in all four tasks of NYUDv2. Notably, MTMamba++ shows significant improvements over MLoRE [14] by +1.05 (mIoU) and +0.97 (odsF) in semantic segmentation and object boundary detection tasks, which demonstrates the superiority of MTMamba++. Moreover, MTMamba++ performs better than MTMamba, showing the effectiveness of S-CTM and LiteHead.
On the PASCAL-Context dataset, MTMamba++ continues to demonstrate superior performance on all tasks except the normal prediction task, which is also comparable. Compared with MLoRE [14], MTMamba++ achieves notable improvements of +0.53 (mIoU), +2.35 (mIoU), +0.66 (maxF), and +3.18 (odsF) in semantic segmentation, human parsing, saliency detection, and object boundary detection tasks, respectively. When compared to the diffusion-based method TaskDiffusion [34], MTMamba++ shows advantages of +0.73 (mIoU), +3.25 (mIoU), +0.62 (maxF), and +3.71 (odsF) in four tasks. These results clearly demonstrate the effectiveness of MTMamba++ for multi-task dense prediction. Furthermore, MTMamba++ outperforms our preliminary work MTMamba on three of five tasks while maintaining comparable performance on the remaining two, further validating the effectiveness of our proposed components.
Table II shows the results on the Cityscapes dataset. We can see that Mamba-based methods perform largely better than the previous CNN-based and Transformer-based approaches on both two tasks. Moreover, MTMamba++ archives the best performance. Notably, MTMamba++ outperforms TaksPrompter [12] by +6.72 (mIoU) in the semantic segmentation task, demonstrating that MTMamba++ is more effective. Besides, MTMamba++ performs better than MTMamba, which shows the effectiveness of S-CTM and LiteHead.
TABLE III: Effectiveness of each core component on NYUDv2. “Multi-task” denotes an MTL model where each task uses standard Swin Transformer blocks [25] after the ECR block in each decoder stage. “Single-task” is the single-task counterpart of “Multi-task”. #11 is the default configuration of MTMamba++.
@@
4 2*Swin LiteHead 53.37 0.5201 19.62 78.40 -0.78 302 436
5 3*Swin DenseHead 54.22 0.5225 19.84 77.40 -1.11 341 563
6 3*Swin LiteHead 54.44 0.5117 19.65 78.60 +0.14 339 533
@@
The qualitative comparisons with baselines (i.e., InvPT [9], TaskPrompter [12], and MTMamba [24]) on NYUDv2, PASCAL-Context, and Cityscapes datasets are shown in Figures 6, 7, and 8, demonstrating that MTMmaba++ provides more precise predictions and details.